Grade 10Physical Chemistry

10gm of substance 80% of it converted into energy so that amount of energy released is ..... mev

Tasneem
7 Years agoGrade 10

1 Answer

Saurabh Koranglekar
7 Years ago
Dear student

80% of 10 gm is 8 gm

so by E= m*c*c = 8 *3*3*10^(8+8-3) j
so E in meV = 72*10^(13) *1.6*10^(-16)
=0.1152

Regards